# Fourier Boundary Features Network with Wider Catchers for Glass Segmentation

### v1

@@ II-B Context Features of Glass
The instability of boundary constraints and weakly blurred boundaries provided by glass in some complex scenes [9] limit the favorable influence of glass boundary detection methods on the accuracy of segmentation results. Therefore, the methods are gradually biased towards the study of how to efficiently leverage abundant context for glass segmentation from multi-scale features extraction focusing [4][10, 11] and effective fusion [12, 13]. Yu et al. [14] have considered deep-layer features with more high-level semantics to be better at the location. Shallow layer features own larger spatial sizes with detailed low-level information instead. PGSNet is designed through progressively aggregating to avoid fusing these features naively. Hu et al. [15] proposed a VIT-based deep architecture to associate multi-layer receptive field features and optimize fusion results Besides, due to the large and complete glass regions context, the attention mechanism is in full utilization. CAGNet [16] integrates two kinds of convolutional attention mechanisms with a Transformer head for deep-shallow feature analysis and fusion. Hou et al. [17] further introduce the Transformer encoder-decoder architecture to boost the usage of multi-levels of context information. Lin et al. [18] embed context correlation into the attention module to bridge contextual links among objects both spatially and semantically.
@@
Figure 2: Architecture of the proposed FBWC. Given an input glass image, there will be two branches to receive the same picture and change them into different features. In particular, one of features is fed into WCC to alleviate the over-capturing and boundary constraints of scarcity. The others is processed by the CTA to supplement and focus fine-grained information. The designed FCC as a learnable feature fusion regulator to balance heterogeneous inputs. Finally, a fine-grained weighting factor regulating the shallow and large area segmentation result with supplementing boundary information constraints to achieve the purpose of dynamic optimized network predictions.
@@ III-A Architecture
Based on the self-learning fusion feature controller, the FBWC derives a lot of adjustment flexibility and self-correction margin. Thus, instead of designing a complex loss function design, we introduce ohem cross-entropy and binary cross-entropy loss to monitor the propagation of semantic information across the proposed network. To be specific, for the $i$th decoder of CU in WCC, binary cross-entropy loss with boundary-awareness is applied by the Canny operator to gradually strengthen boundary constraints. At the same time, $\ell_{seg}$ and $\ell_{aux}$ refer to supervised learning from ground truth to the controller and attention mechanism results with ohem cross-entropy loss , respectively. Therefore, the final loss for FBWC is:
$Loss=\ell_{seg}+\ell_{aux}+\frac{1}{I}\sum_{i=1}^{I}\ell_{i}.$ (1)
@@ III-B Wider Coarse-Catchers
As shown in Figure 3, we employ CUs as the component of our WCC. We notice that shallow encoded information can provide a large area of segmentation, and the decoding process is conducive to boundary constraints. Hence, the CU consists of three layers $l\in[1,2,3]$ encoding-decoding symmetric structure and three important points: start point, trough point, and end point. In particular, the constraint point is named by sharing the both start and end position in WCC.
Specifically, in the encoding process, we use convolution and down-sampling to double the number of channels and half the size of height and width respectively on each layer, which can be formulated as:
$x_{En,l}^{i}=\Gamma^{l-1}\left\{\omega\cdot\vartheta^{3\times 3}\left(x_{En,1}^{i}\right)\right\},l\in[2,3],$ (2)
where $\omega$ and $\vartheta^{3\times 3}$ denote dowm-sampling and convolutional operator, respectively. And the $\Gamma^{l}\{\cdot\}$ is defined as the continuous operator which executes $l$ times.
For the $i$th CU, $x_{En,3}^{i}$ or $x_{De,3}^{i}$ and $x_{En,1}^{i}$ are named trough point and start point, respectively. After the continuous operation of convolution and dowmsampling, in then process of decoding, we apply up-sampling operators with bilinear interpolation $\mu$ and the skip connection $S(\cdot)$ to bridge result features of up-sampling and $x_{En,l}^{i}$. Finally, the decoding process is defined as:
$x_{De,l}^{i}=\vartheta^{3\times 3}\left[S\left(\mu\left(x_{De,l+1}^{i}\right),x_{En,l}^{i}\right)\right],l\in[1,2],$ (3)
where $x_{De,l}^{i}$ denotes the features of the $l$th layer on the decoder of $i$th CU. And the operator of skip connection $S(\cdot)$ is defined as $S(a,b)=a\oplus b$. To sum up, the CU mapping ( $\Xi_{CU}$ ) can be expressed as:
$\Xi_{CU}(z)=S(\check{z},\tilde{z}),$ (4)
@@
where ${\check{z}}$ and ${\hat{z}}$ represent continuously downsampling-convolutional and upsampling-convolutional results based on the feature $z$, respectively. Furthermore, $\tilde{z}$ is denoted as upsampling the results of ${\check{z}}$. The process of input $x_{\text{cir }}$ getting outputs after the first CU is formulated as:
@@
$x_{En,3}^{1}$ is for coarse-grained features acquisition and $x_{De,1}^{1}$ is for continuous asymptotic boundary constraints. Generally, when $i\in[1,2,\ldots,N]$, this means that multiple CUs need to be connected in series. As a result, the representation of variables on output layers in CUs are defined as following:
$\left\{\begin{array}[]{l}x_{En,1}^{1}=\omega\cdot\vartheta^{3\times 3}\left(x_{cir}\right),i=1\\ x_{En,1}^{i}=x_{De,1}^{i-1},i\neq 1.\\ \left[\begin{array}[]{l}x_{De,1}^{i}\\ x_{En,3}^{i}\end{array}\right]=\Gamma^{i}\left\{\Xi_{CU}\left(x_{cir}\right)\right\},others\end{array}\right.$ (6)
Finally, the constraint points set $X_{De,1}=\left\{x_{De,1}^{i}\right\}_{i=1}^{N}$ and trough points set $X_{En,3}=\left\{x_{En,3}^{i}\right\}_{i=1}^{N}$ are represented as the outputs of multiple CUs for forming a wider shallow feature capture framework.
@@ III-C Cross Transpose Attention
where $\sigma$ refers to Sigmoid function and $CTA(a,b)$ is denoted as Cross Transpose Attention operator with tensors of $a$ and $b$. The cross similarity coefficients are expressed by $\alpha$ and $\beta$. The final output of CTA is $x_{tri}^{\text{out }}=CTA\left(\left(CTA\left(x_{tri},x_{En,1}^{cir}\right),x_{En,2}^{cir}\right)\right)$ for effective focusing on fine-grained semantic features.
@@ III-D Fourier Convolution Controller
$X^{F}=\vartheta^{1\times 1}(X_{En,3})+{\boxplus}_{x^{i}_{En,3}}[F_{x_{En,3}}(c,d)],i\in[1,2,3,4],$ (11)
where the ${\boxplus}_{x^{i}_{En,3}}[\cdot]$ means executing operation for each channel of $x^{i}_{En,3}$ which belongs to $X_{En,3}$, and the $F_{x_{En,3}}(c,d)$ represents the glass edge information extracting process:
$F_{x_{En,3}}(c,d)=\sum_{w=0}^{\frac{W}{\lambda}-1}\sum_{h=0}^{\frac{H}{\lambda}-1}x_{En,3}(w,h)e^{-j2\pi\left(\frac{\lambda cw}{W}+\frac{\lambda dh}{H}\right)},$ (12)
where $c\in[0,\ \dots\ ,\frac{W}{\lambda}-1]$ and $d\in[0,\ \dots\ ,\frac{H}{\lambda}-1]$ refer to frequency domain variables. In the FFT process, the $x_{En,3}(w,h)$ is the value of the feature point at position $(w,h)$ in one channel belongs to $x^{i}_{En,3}$.
Proposition 1. Assuming $I^{*}$ is an image with regular geometric boundaries. For any continuous piecewise linear function subset $\Omega$ and any $f\in\Omega$, any convolutional neural network $f_{J}^{\mathbf{w},\mathbf{b}}$ with depth $J\in\mathbb{Z^{+}}$ and ReLU function exists:
$\left\|f-f_{J}^{\mathbf{w},\mathbf{b}}(FFT(I^{*}).real+I^{*})\right\|_{\Omega}<\left\|f-f_{J}^{\mathbf{w},\mathbf{b}}(I^{*})\right\|_{\Omega}.$ (13)
Proof. The two-dimensional FFT can be regarded as a generalization of one-dimensional FFT, then we obtain expression of a single feature point value on a single column:
$x_{En,3}(h)=\frac{a_{0}}{2}+\sum_{n=1}^{P}\left[a_{n}\cdot\cos(nwh)+b_{n}\cdot\sin(nwh)\right].$ (14)
We replace $\cos(nwh)$ and $\sin(nwh)$ by $\left(e^{inwh}+e^{-inwh}\right)/2$ and $-i\left(e^{inwh}-e^{-inwh}\right)/2$ respectively, then Eq.14 can be reformed as:
$x_{En,3}(h)=\frac{a_{0}}{2}+\sum_{n=1}^{P}\left[\frac{a_{n}-ib_{n}}{2}\cdot e^{inwh}+\frac{a_{n}+ib_{n}}{2}\cdot e^{-inwh}\right].$ (15)
@@
$x_{En,3}(h)$ $=c_{0}\cdot e^{ihw\cdot 0}+c_{1}\cdot e^{ihw\cdot 1}+c_{2}\cdot e^{ihw\cdot 2}+\ \dots$ (16)
$+c_{N-1}\cdot e^{ihw\cdot N-1},w=2\pi/N.$
@@
$\Theta\cdot\left[\begin{array}[]{c}c_{0}\\ c_{1}\\ c_{2}\\ \vdots\\ c_{N-1}\end{array}\right]=\left[\begin{array}[]{c}x_{En,3}(0)\\ x_{En,3}(1)\\ x_{En,3}(2)\\ \vdots\\ x_{En,3}(N-1)\end{array}\right],$ (17)
@@
Then we use $\Theta^{-1}(j,k)$ to represnet the element in $j$th row and $k$th column in $\Theta^{-1}$, then $c_{j}$ can be expressed as:
@@
Figure 5: Illustration of FCC, which serves as the core component of FBWC for providing a flexible solution to the features fusion problem.
@@
then we transfer Eq.19 into the following form:
@@
From Eq.11 and Eq.23, the FFT result is shown as follow:
$X^{F}=\vartheta^{1\times 1}(X_{En,3})+{\boxplus}_{x^{i}_{En,3}}[\sum_{w=0}^{\frac{W}{\lambda}-1}\sum_{l=1}^{\log_{2}N}x_{En,3(l)}]$ (24)
$,i\in[1,2,3,4],l\in[1,2,\ldots,\log_{2}N].$

### v2

@@ II-B Context Features of Glass
The instability of boundary constraints and weakly blurred boundaries provided by glass in some complex scenes [9] limit the favorable influence of glass boundary detection methods on the accuracy of segmentation results. Therefore, the methods are gradually biased towards the study of how to efficiently leverage abundant context for glass segmentation from multi-scale features extraction focusing [4][10, 11] and effective fusion [12, 13]. Yu et al. [14] have considered deep-layer features with more high-level semantics to be better at the location. Shallow layer features own larger spatial sizes with detailed low-level information instead. PGSNet is designed through progressively aggregating to avoid fusing these features naively. Hu et al. [15] proposed a VIT-based deep architecture to associate multi-layer receptive field features and optimize fusion results. Besides, due to the large and complete glass regions context, the attention mechanism is in full utilization. CAGNet [16] integrates two kinds of convolutional attention mechanisms with a Transformer head for deep-shallow feature analysis and fusion. Hou et al. [17] further introduce the Transformer encoder-decoder architecture to boost the usage of multi-levels of context information. Lin et al. [18] embed context correlation into the attention module to bridge contextual links among objects both spatially and semantically.
@@
Figure 2: Architecture of the proposed FBWC. Given an input glass image, there will be two branches to receive the same picture and change them into different features. In particular, one of features is fed into WCC to alleviate the over-capturing and boundary constraints of scarcity. The others is processed by the CTA to supplement and focus fine-grained information. The designed FCC as a learnable feature fusion regulator to balance heterogeneous inputs. Finally, a fine-grained weighting factor regulating the shallow and large area segmentation result with supplementing boundary information constraints to achieve the purpose of dynamic optimized network predictions. The number next to each module denotes its parameter. Express model concerns more intuitively by exporting Hot Map.
@@ III-A Architecture
Based on the self-learning fusion feature controller, the FBWC derives a lot of adjustment flexibility and self-correction margin. Thus, instead of designing a complex loss function design, we introduce ohem cross-entropy [28] and binary cross-entropy loss to monitor the propagation of semantic information across the proposed network. To be specific, for the $i$-th decoder of CU in WCC, binary cross-entropy loss with boundary-awareness is applied by the Canny operator to gradually strengthen boundary constraints. At the same time, $\ell_{Seg}$ and $\ell_{AM}$ refer to supervised learning from ground truth to the controller and Attention Mechanism (AM) results with ohem cross-entropy loss , respectively. Therefore, the final loss for FBWC is:
$Loss=\ell_{Seg}+\ell_{AM}+\frac{1}{{N_{CUs}}}\sum_{{\overline{i}}=1}^{N_{CUs}}\ell_{\overline{i}},$ (1)
where ${N_{CUs}}$ represents numbers of CUs, and $\{\ell_{\overline{i}}|{\overline{i}}\in{N_{CUs}}\}$ denotes the loss of Boundary Constraint (BC Loss) to a single CU.
@@ III-B Wider Coarse-Catchers
As shown in Figure 3, we employ CUs as the component of our WCC. We notice that shallow encoded information can provide a large area of segmentation, and the decoding process is conducive to boundary constraints. Hence, the CU consists of three layers $\overline{l}\in[1,2,3]$ encoding-decoding symmetric structure and three important points: start point, trough point, and end point. In particular, the constraint point is named by sharing the both start and end position in WCC.
Specifically, in the encoding process, we use convolution and down-sampling to manage the number of channels and half the size of height and width respectively on each layer, which can be formulated as:
$x_{En,\overline{l}}^{\overline{i}}=\Gamma^{\overline{l}-1}\left\{\omega\cdot\vartheta^{3\times 3}\left(x_{En,1}^{\overline{i}}\right)\right\},\overline{l}\in[2,3],$ (2)
where $\omega$ and $\vartheta^{3\times 3}$ denote dowm-sampling and convolutional operator, respectively. And the $\Gamma^{{\overline{l}}}\{\cdot\}$ is defined as the continuous operator which executes ${\overline{l}}$ times.
For the ${\overline{i}}$-th CU, $x_{En,3}^{\overline{i}}$ or $x_{De,3}^{\overline{i}}$ and $x_{En,1}^{\overline{i}}$ are named trough point and start point, respectively. After the continuous operation of convolution and dowmsampling, in the process of decoding, we apply up-sampling operators with bilinear interpolation $\mu$ and the skip connection to bridge result features of up-sampling and $x_{En,{\overline{l}}}^{\overline{i}}$. Finally, the decoding process is defined as:
$x_{De,{\overline{l}}}^{\overline{i}}=\vartheta^{3\times 3}\left[Skip\left(\mu\left(x_{De,{\overline{l}}+1}^{\overline{i}}\right),x_{En,{\overline{l}}}^{\overline{i}}\right)\right],{\overline{l}}\in[1,2],$ (3)
where $x_{De,{\overline{l}}}^{\overline{i}}$ denotes the features of the ${\overline{l}}$-th layer on the decoder of ${\overline{i}}$-th CU. And the operator of skip connection $Skip(\cdot,\cdot)$ is defined as $Skip(a,b)=a\oplus b$. To sum up, the CU mapping ( $\Xi_{CU}$ ) can be expressed as:
$\Xi_{CU}(z)=Skip(\check{z},\tilde{z}),$ (4)
@@
where ${\check{z}}$ represents continuously downsampling via convolution. Furthermore, $\tilde{z}$ is denoted as upsampling the results of ${\check{z}}$. The process of input $x_{\text{cir }}$ getting outputs after the first CU is formulated as:
@@
where $x_{En,3}^{1}$ is for coarse-grained features acquisition and $x_{De,1}^{1}$ is for continuous asymptotic boundary constraints. Generally, when ${\overline{i}}\in[1,2,\ldots,N]$, it means that multiple CUs need to be connected in series. As a result, the representation of variables on output layers in CUs are defined as following:
$\left\{\begin{array}[]{l}x_{En,1}^{1}=\omega\cdot\vartheta^{3\times 3}\left(x_{cir}\right),{\overline{i}}=1\\ x_{En,1}^{\overline{i}}=x_{De,1}^{{\overline{i}}-1},{\overline{i}}\neq 1.\\ \left[\begin{array}[]{l}x_{De,1}^{\overline{i}}\\ x_{En,3}^{\overline{i}}\end{array}\right]=\Gamma^{\overline{i}}\left\{\Xi_{CU}\left(x_{cir}\right)\right\},others\end{array}\right.$ (6)
Finally, the constraint points set $X_{De,1}=\left\{x_{De,1}^{\overline{i}}\right\}_{{\overline{i}}=1}^{N}$ and trough points set $X_{En,3}=\left\{x_{En,3}^{\overline{i}}\right\}_{{\overline{i}}=1}^{N}$ are represented as the outputs of multiple CUs for forming a wider shallow feature capture framework.
@@ III-C Cross Transpose Attention
where $\sigma$ refers to Sigmoid function and $CTA(a,b)$ is denoted as Cross Transpose Attention operator with tensors $a$ and $b$. The cross similarity coefficients are expressed by $\alpha$ and $\beta$. The final output of CTA is $x_{tri}^{\text{out }}=CTA\left(\left(CTA\left(x_{tri},x_{En,1}^{cir}\right),x_{En,2}^{cir}\right)\right)$ for effective focusing on fine-grained semantic features.
@@ III-D Fourier Convolution Controller
We use $X^{F}$ to represent the result of $\underset{\varpi}{g_{2}}\left(X_{En,3}\right)$ which fused the glass frequency domain information obtained through FFT:
$X^{F}=\vartheta^{1\times 1}(X_{En,3})+{\boxplus}_{x^{\overline{i}}_{En,3}}[F_{x_{En,3}}(c,d)],{\overline{i}}\in[1,2,3,4],$ (11)
where the ${\boxplus}_{x^{\overline{i}}_{En,3}}[\cdot]$ means executing operation for each channel of $x^{\overline{i}}_{En,3}$ which belongs to $X_{En,3}$, and the $F_{x_{En,3}}(c,d)$ represents the glass frequency domain information extracting process:
$F_{x_{En,3}}(c,d)=\sum_{w^{*}=0}^{\frac{W}{\lambda}-1}\sum_{h^{*}=0}^{\frac{H}{\lambda}-1}x_{En,3}(w^{*},h^{*})e^{-j2\pi\left(\frac{\lambda cw^{*}}{W}+\frac{\lambda dh^{*}}{H}\right)},$ (12)
where $c\in[0,\ \dots\ ,\frac{W}{\lambda}-1]$ and $d\in[0,\ \dots\ ,\frac{H}{\lambda}-1]$ refer to frequency domain variables. In the FFT process, the $x_{En,3}(w^{*},h^{*})$ is the value of the feature point at position $(w^{*},h^{*})$ in one channel belongs to $x^{\overline{i}}_{En,3}$.
The two-dimensional FFT can be regarded as a generalization of one-dimensional FFT, then we obtain expression of a single feature point value on a single column:
$x_{En,3}(\cdot,h^{*})=\frac{a_{0}}{2}+\sum_{n=1}^{P}\left[a_{n}\cdot\cos(nwh^{*})+b_{n}\cdot\sin(nwh^{*})\right].$ (13)
We replace $\cos(nwh^{*})$ and $\sin(nwh^{*})$ by $\left(e^{inwh^{*}}+e^{-inwh^{*}}\right)/2$ and $-i\left(e^{inwh^{*}}-e^{-inwh^{*}}\right)/2$ respectively, then Formula (13) can be reformed as:
$x_{En,3}(\cdot,h^{*})=\frac{a_{0}}{2}$ (14)
$+\sum_{n=1}^{P}\left[\frac{a_{n}-ib_{n}}{2}\cdot e^{inwh^{*}}+\frac{a_{n}+ib_{n}}{2}\cdot e^{-inwh^{*}}\right].$
@@
$x_{En,3}(\cdot,h^{*})$ $=c_{0}\cdot e^{ih^{*}w\cdot 0}+c_{1}\cdot e^{ih^{*}w\cdot 1}+c_{2}\cdot e^{ih^{*}w\cdot 2}+\ \dots$ (15)
$+c_{N-1}\cdot e^{ih^{*}w\cdot N-1},w=2\pi/N.$
@@
$\Theta\cdot\left[\begin{array}[]{c}c_{0}\\ c_{1}\\ c_{2}\\ \vdots\\ c_{N-1}\end{array}\right]=\left[\begin{array}[]{c}x_{En,3}(\cdot,0)\\ x_{En,3}(\cdot,1)\\ x_{En,3}(\cdot,2)\\ \vdots\\ x_{En,3}(\cdot,N-1)\end{array}\right],$ (16)
@@
Then we use $\Theta^{-1}(j,k)$ to represnet the element in $j$-th row and $k$-th column in $\Theta^{-1}$, then $c_{j}$ can be expressed as:
@@
Figure 5: Illustration of FCC, which serves as the core component of FBWC for providing a flexible solution to the features fusion problem. Express model concerns more intuitively by exporting Hot Map.
@@
Then we transfer Formula (19) into the following form:
@@
From Formula (11) and Formula (22), the FFT result is shown as follow:
$X^{F}=\vartheta^{1\times 1}(X_{En,3})+{\boxplus}_{x^{\overline{i}}_{En,3}}[\sum_{w^{*}=0}^{\frac{W}{\lambda}-1}\sum_{l=1}^{\log_{2}N}x_{En,3(l)}(w^{*})]$ (23)
$,{\overline{i}}\in[1,2,3,4],l\in[1,2,\ldots,\log_{2}N].$